OpenStudy (anonymous):

find derivative of f(x)=x(1-(4/x+3))^1/4

OpenStudy (anonymous):

\[f(x)=x(1-\frac{ 4 }{ x+3})^{\frac{ 1 }{ 4 }}\] is it like this??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you know how to use the product rule?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so you need to use that rule for this, and tell me what you got:) if you have any difficulty just ask me...

OpenStudy (anonymous):